# A Comprehensive Study of Shapley Value in Data Analytics [Experiment, Analysis & Benchmark]

### v6

@@ 5. Evaluation
(c) The approximation error measured by $\epsilon=1-cosine\_similarty(\vec{\hat{\phi}},\vec{\phi})$ (Liu et al., 2022a).
@@
Figure 5. Impacts of approximation error (x-axis: the approximation error $\epsilon$).
In this section, we use SVBench to implement multiple SV computing algorithms, including both the base algorithms that have been studied by prior works and hybrid algorithms with novel combinations of SV computing techniques. Using the implemented algorithms, we not only validate the usability, modularity, and flexibility of SVBench but also study the following four sets of evaluations in order to answer the aforementioned questions:
  • §5.4 explores the relationship between SV of the four types of players in DA and the overall utility of their associated tasks, offering insights to the question bold in §3.2.4.
Table 9 summarizes detailed configurations specified for each algorithm. We note that this work is the first attempt to combine the $\spadesuit$-tagged base algorithm with the five optimization techniques. We use MLE as the base algorithm in §5.2–§5.4, since it generally achieves better efficiency and accuracy performance than the other base algorithms in §5.1 and varying base algorithms would not influence conclusions in those subsections.
For the generality of findings from experiments, we select 10 canonical datasets from a large number of literature to conduct four types of DA tasks – result interpretation (RI), data tuple valuation (DV), dataset valuation (DSV), and federated learning (FL), which enable evaluations of SV for different types of players defined in the current DA domain. Table 10 summarizes the datasets, including Iris (Markelle Kelly, [n.d.]), Wine (Markelle Kelly, [n.d.]), Mnist (LeCun and Cortes, 2010), Cifar-10 (Krizhevsky, 2009), Adult (Markelle Kelly, [n.d.]), Tic-Tac-Toe (Ttt) (Markelle Kelly, [n.d.]), Bank (Markelle Kelly, [n.d.]), Dota2 (Markelle Kelly, [n.d.]), Wind (Vanschoren et al., 2013; Feurer et al., 2020), 2Dplanes (Vanschoren et al., 2013; Feurer et al., 2020), and our task settings. More details of task settings and the code are available at GitHub [^2] https://github.com/zjuDBSystems/SVBench. .
5.1. Computation Efficiency
This section investigates the efficiency performance of five base SV computing algorithms (MC, RE, MLE, GT, CP) and the hybrid algorithms, each of which selects an optimization strategy from Table 8 and integrates this strategy with a base algorithm. To control the time cost of each experiment, we follow previous work (Liu et al., 2022a) to monitor the approximation stability by $\Delta\hat{\phi}=\frac{1}{5n}\sum_{m=1}^{5}\sum_{i=1}^{n}|\frac{\hat{\phi}^{e}_{i}-\hat{\phi}^{e-m\times n}_{i}}{\hat{\phi}^{e}_{i}}|$ and terminate the approximation when the convergence criterion ($\Delta\hat{\phi}<\tau$) is satisfied, where $\hat{\phi}^{e}_{i}$ is the approximate SV of player $p_{i}$ after $e$ times of utility computation, $\tau$ is the convergence threshold set to 0.05 for all tasks. We measure the efficiency performance by total time cost $N_{uc}\times T_{uc}$, and show the computation complexity $N_{uc}$ and the approximation error $\epsilon$, as a reference in the results.
As denoted by blue bars, MLE generally achieves the leading efficiency performance (lower $T_{uc}\times N_{uc}$ and $N_{uc}$) among the five base algorithms. The major reason is that, unlike MLE, algorithms such as RE, GT, and CP must solve auxiliary optimization problems (e.g., weighted least squares in RE, feasibility constraints in GT, or convex objectives in CP) in the computation, introducing extra cost. In particular, when it is hard to solve the corresponding optimization problem for the DA task (e.g., RI-Adult, RI-2Dplanes, and DV-Wind), prolonged convergence will appear. Although CP can converge faster than MLE in some tasks, e.g., RI-Iris, RI-Ttt, DV-Iris, it has strong dependency on the sparse SV assumption and performs well only when that assumption is satisfied (to some extent). We also notice that MC can outperform MLE occasionally (e.g., in DSV-Cifar, FL-Dota2). However, MLE generally achieves higher accuracy (lower $\epsilon$) than MC under identical convergence criteria, due to its closed-form integral expression of SV, enabling more deterministic computation with high precision (e.g., via Gaussian quadrature).
Research Direction 1: Exploration on innovative hybrid SV computing algorithms. Our evaluations show that the hybrid SV computing algorithms integrating multiple efficiency optimization techniques achieve better performance than the algorithms using only one of the integrated techniques in most cases. We highly recommend combining TC with a base SV computing algorithm in all DA tasks and activating GA and TSS when the utility computation (e.g., in FL) needs costly model training and evaluation. For the selection of base algorithms, it is recommended to choose MLE (or MC). In summary, innovative hybrid SV computing algorithms with extensive experiments are anticipated for efficient SV computation in the literature.
@@ 5.1. Computation Efficiency
5.2. Approximation Error
Figure 5(a) presents the relationship between the approximation error of three SV computing algorithms and the computation efficiency and stability, while Figure 5(b) presents the impacts of approximation error on the effectiveness of SV. The square-tagged lines in Figure 5(a) show that, in most tasks, the larger the SV approximation error, the smaller the computation complexity. This is due to far fewer utility values needing to be computed when a larger tolerance is given to the approximation error, regardless of the strategy for sampling coalitions as discussed in §3.2.2. However, as shown in the star-tagged lines in Figure 5(a), the approximate SVs with a large error tend to be generated when the computation is far from an ideal convergence status (where $\Delta\hat{\phi}$ goes below a trivial threshold value). Those SVs may lead to a larger error in the scaled SV results, as shown in Figure 5(b), indicating the potential ineffectiveness of using approximated SV for pricing, selection, weighting, and attribution in DA. We notice that the RI_Iris task’s results in Figure 5(a) show a different pattern from the other tasks. This is for the same reason as the performance of this task in §5.1, that only four players in this task make all algorithms complete full sampling before converging. In conclusion, the above results consolidate the necessity to strike a balance among the approximation error, computation complexity, and the effectiveness of SV. To achieve such a balance, we highlight the following future directions.
Research Direction 2: Investigation on runtime dynamic tuning of convergence criterion when using sampling-based approximation techniques. Take the criterion $\Delta\hat{\phi}<\tau$ and the DV task in Figure 1 as an example. A small threshold (e.g.,$\tau=10^{-5}$) can be set at the start of SV approximation. Then, the task can periodically check whether image samples ranked in the top 50% based on the latest approximated SV produce more accurate classification models than those produced by images ranked in the bottom. Once satisfied, the approximation can be terminated to save computation cost.
@@ 5.2. Approximation Error
5.3. Privacy Protection
This section studies the effectiveness of three popular privacy protection techniques (DP, QT, DR) for preventing SV-driven attacks, including FIA (Luo et al., 2022) which may occur in RI tasks, and MIA (Wang et al., 2023) which may occur in DV task. We also study the impacts of the three techniques on SV’s effectiveness.
The implementation of two attacks and three countermeasures is based on previous papers (Luo et al., 2022; Wang et al., 2023). We follow Luo et al. (2022) to use two random distributions, Uniform(0,1) and Gaussian(0.5,$0.25^{2}$), to perform FIA (called FIAU and FIAG respectively), and follow Wang et al. (2023) to implement MIA. We tune the strength of privacy protection from low to high by varying (1) the standard deviation of noise generated by DP from 0.1, 0.5 to 0.9, (2) the number of distinctive discrete Shapley values produced by QT from $0.9n$, $0.5n$ to $0.1n$, and (3) the number of dimensions reduced for $\vec{\phi}$ by DR from $0.1n$, $0.5n$ to $0.9n$. We measure the performance in preventing SV-driven FIA by the MAE metric used by Luo et al. (2022) and measure the effectiveness of preventing SV-driven MIA by the AUROC score used by Wang et al. (2023). As both RI and DV tasks rely on the ranking of the final SV to find the top important data features and samples, the impact of privacy protection techniques on SV effectiveness is measured by the variance in SV ranking results before and after protection.
For preventing FIA, DP is the most effective, while QT and DR need tuning the privacy protection strength to a high level for an apparent MAE enlargement. Despite this, DP renders a much larger impact on SV ranking (thus affects more on SV’s effectiveness) under the setting of low or median privacy protection strength. These phenomena are mainly because DP adds noise to SV of all features no matter how the privacy protection strength is varied, while QT and DR alter SV of only 10% (or 50%) of data features when the strength is set to a low (or median) level.
For preventing MIA, QT and DR are effective in most cases, while DP leads to AUROC larger than the results achieved without privacy protection in some cases, e.g., DV-Bank, DV-Ttt. Moreover, different from the results of preventing FIA, enhancing the privacy protection strength may not result in better effectiveness (lower AUROC) of the three techniques in preventing MIA. For example, in DV-Wind, QT performs the worst when setting a high privacy protection strength. These outcomes stem from the inability of QT, DR, and DP to rigorously enforce indistinguishability between two Shapley value distributions: the IN distribution, computed when the target sample is included in the dataset, and the OUT distribution, computed when the target sample is excluded (Wang et al., 2023). Specifically, due to randomness, DP may inject markedly dissimilar noise into SVs used for generating IN and OUT distributions and thus enlarge the difference between the two distributions, making the success of MIA much easier. QT and DR map the original SVs into a new value space, in which the difference between SVs used for generating IN and OUT distributions can be more distinguishable, negatively affecting the efficacy in preventing SV-driven MIA.
@@ 5.4. SV Interpretations
This section examines the correctness of the mainstream SV interpretation paradigm, i.e., utility-based, in explaining SV of different types of players defined in the current DA domain. This paradigm points out that the larger the SV of a player, the more the impacts that the player could pose on the DA task’s overall utility. Therefore, we observe the change in the overall utility of DA tasks caused by removing or adding players.
@@
Research Direction 6: In-depth investigation on key factors affecting SV interpretations. SV may not be correctly interpreted or fail to serve its desired application purposes due to two points. One is the difference between the exact SV and the approximate SV. The other one is the mismatch between the DA task and the implicit assumption of SV. For example, DV-Wind and DSV-2Dplanes tasks as discussed above, mismatching with the assumption that the average contribution is appropriate to fairly distribute the overall utility of the task among players. Another example is the RI task in Figure 1, where pixels in an image interact mutually to influence the model predictions on the label of this image, mismatching with the assumption that the contribution of any individual player to the cooperative game is independent of its interactions with other players. The work mining the deep understanding of these factors under more types of DA tasks to enhance the effectiveness of SV is anticipated.
Research Direction 7: Exploration on supports for complicated DA tasks. The real-world DA tasks are complicated, containing mutually dependent players (Chu and Chan, 2020; Olsen et al., 2022; Aas et al., 2021b, a; Frye et al., 2021) and real-time dynamic updates on the player set (Sun et al., 2024; Xia et al., 2025). For example, the RI task in Figure 1, where the influence of one pixel on the prediction from the image classification model tends to interrelate with nearby pixels. Besides, the dynamic real-time updates of the player set (Sun et al., 2024; Xia et al., 2025) are common in analytical tasks built upon the databases (Zhang et al., 2023b; Xia et al., 2025; Padala et al., 2025). However, most existing SV applications are ill-suited for these tasks due to the assumptions that the players are independent and the player set is static and immutable. Although conditional SV (Chen et al., 2023) and dynamic SV (Zhang et al., 2023b) have been devised for complicated tasks, these new types of SV do not treat the mutual dependency among players and their real-time dynamic updates as a single interrelated problem, thus limiting in practical applications. It is expected that future work can devise more SV applications under the complicated yet realistic DA task settings.
@@ 6. Conclusion
This paper comprehensively studied the Shapley value applied throughout the data analytics workflow. We summarized the critical variables (i.e., the player and the utility function) in designing SV applications for DA and clarified the essential functionalities of SV for data scientists. We condensed the technical challenges of applying SV in DA and discussed the related arts, qualitatively and quantitatively. The conclusion of experimental evaluations based on our development framework, SVBench, supports our findings in a synthetic review. At last, we identified the limitations of the current efforts and offered insights into the directions of future work.

### v7

@@ 5. Evaluation
(c) The approximation error measured by $\epsilon$.
@@
In this section, we use SVBench to implement multiple SV computing algorithms, including both the base algorithms that have been studied by prior works and the hybrid algorithms with novel combinations of SV computing techniques. Using the implemented algorithms, we not only validate the usability, modularity, and flexibility of SVBench but also study the following four sets of evaluations in order to answer the aforementioned questions:
  • §5.4 explores the relationship between the SVs of the four types of players in DA and the overall utility of their associated tasks, offering insights to the question bold in §3.2.4.
Table 9 summarizes the detailed configurations specified for each algorithm. We note that this work is the first attempt to combine the $\spadesuit$-tagged base algorithm with the five optimization techniques. We use MLE as the base algorithm in §5.2–§5.4, since it generally achieves better efficiency and accuracy performance than the other base algorithms in §5.1 and varying base algorithms would not influence conclusions in those subsections.
For the generality of findings from experiments, we select 10 canonical datasets from a large number of literature to conduct four types of DA tasks – result interpretation (RI), data tuple valuation (DV), dataset valuation (DSV), and federated learning (FL), which enable evaluations of SV for different types of players defined in the current DA domain. Table 10 summarizes the datasets, including Iris (Markelle Kelly, 2025), Wine (Markelle Kelly, 2025), Mnist (LeCun and Cortes, 2010), Cifar-10 (Krizhevsky, 2009), Adult (Markelle Kelly, 2025), Tic-Tac-Toe (Ttt) (Markelle Kelly, 2025), Bank (Markelle Kelly, 2025), Dota2 (Markelle Kelly, 2025), Wind (Vanschoren et al., 2013; Feurer et al., 2020), 2Dplanes (Vanschoren et al., 2013; Feurer et al., 2020), and our task settings. More details of task settings and the code are available at GitHub [^2] https://github.com/zjuDBSystems/SVBench. .
5.1. Computation Efficiency
This section investigates the efficiency performance of five base SV computing algorithms (MC, RE, MLE, GT, CP) and the hybrid algorithms, each of which selects an optimization strategy from Table 8 and integrates this strategy with a base algorithm. To control the time cost of each experiment, we follow previous work (Liu et al., 2022a) to monitor the approximation stability by $\Delta\hat{\phi}=\frac{1}{5n}\sum_{m=1}^{5}\sum_{i=1}^{n}|\frac{\hat{\phi}^{e}_{i}-\hat{\phi}^{e-m\times n}_{i}}{\hat{\phi}^{e}_{i}}|$ and terminate the approximation when the convergence criterion ($\Delta\hat{\phi}<\tau$) is satisfied, where $\hat{\phi}^{e}_{i}$ is the approximate SV of player $p_{i}$ after $e$ times of utility computation, $\tau$ is the convergence threshold set to 0.05 for all tasks. We measure the efficiency performance by the total time cost $N_{uc}\times T_{uc}$, and show the computation complexity $N_{uc}$ and the approximation error $\epsilon=1-\frac{\sum_{i=1}^{n}\hat{\phi}_{i}\cdot\phi_{i}}{\sqrt{\sum_{i=1}^{n}\hat{\phi}_{i}^{2}}\cdot\sqrt{\sum_{i=1}^{n}\phi_{i}^{2}}}$ (Liu et al., 2022a), as a reference in the results.
As denoted by blue bars, MLE generally achieves the leading efficiency performance (lower $T_{uc}\times N_{uc}$ and $N_{uc}$) among the five base algorithms. The major reason is that, unlike MLE, algorithms such as RE, GT, and CP must solve auxiliary optimization problems (e.g., weighted least squares in RE, feasibility constraints in GT, or convex objectives in CP) in the computation, introducing an extra cost. In particular, when it is hard to solve the corresponding optimization problem for the DA task (e.g., RI-Adult, RI-2Dplanes, and DV-Ttt), prolonged convergence will appear. Although CP can converge faster than MLE in some tasks, e.g., RI-Iris, RI-Ttt, DV-Iris, it has a strong dependency on the sparse SV assumption and performs well only when that assumption is satisfied (to some extent). We also notice that MC can outperform MLE occasionally (e.g., in DSV-Cifar, FL-Dota2). However, MLE generally achieves higher accuracy (lower $\epsilon$) than MC under identical convergence criteria, due to its closed-form integral expression of SV, enabling more deterministic computation with high precision (e.g., via Gaussian quadrature).
Research Direction 1: Exploration on innovative hybrid SV computing algorithms. Our evaluations show that the hybrid SV computing algorithms integrating multiple efficiency optimization techniques achieve better performance than the algorithms using only one of the integrated techniques in most cases. We highly recommend combining TC with a base SV computing algorithm in all DA tasks and activating GA and TSS when the utility computation (e.g., in FL) needs costly model training and evaluation. For the base algorithm, it is recommended to choose MLE (or MC). In summary, innovative hybrid SV computing algorithms with extensive experiments are anticipated for efficient SV computation in the literature.
5.2. Approximation Error
@@
Figure 5. Impacts of approximation error (x-axis: the $\epsilon$ value).
Figure 5(a) presents the relationship between the approximation error of three SV computing algorithms and the computation efficiency and stability, while Figure 5(b) presents the impacts of approximation error on the effectiveness of SV. The square-tagged lines in Figure 5(a) show that, in most tasks, the larger the SV approximation error, the smaller the computation complexity. This is due to far fewer utility values needing to be computed when a larger tolerance is given to the approximation error, regardless of the strategy for sampling coalitions, as discussed in §3.2.2. However, as shown in the star-tagged lines in Figure 5(a), the approximate SVs with a large error tend to be generated when the computation is far from an ideal convergence status (where $\Delta\hat{\phi}$ goes below a trivial threshold value). Those SVs may lead to a larger error in the scaled SV results, as shown in Figure 5(b), indicating the potential ineffectiveness of using approximate SV for pricing, selection, weighting, and attribution in DA. We notice that the RI_Iris task’s results in Figure 5(a) show a different pattern from the other tasks. This is for the same reason as the performance of this task in §5.1, that only four players in this task make all algorithms complete full sampling before converging. In conclusion, the above results consolidate the necessity to strike a balance among the approximation error, computation complexity, and the effectiveness of SV. To achieve such a balance, we highlight the following future directions.
Research Direction 2: Investigation on the runtime dynamic tuning of convergence criterion when using sampling-based approximation techniques. Take the criterion $\Delta\hat{\phi}<\tau$ and the DV task in Figure 1 as an example. A small threshold (e.g.,$\tau=10^{-5}$) can be set at the start of SV approximation. Then, the task can periodically check whether image samples ranked in the top 50% based on the latest approximate SV produce more accurate classification models than those produced by images ranked in the bottom. Once satisfied, the approximation can be terminated to save computation cost.
@@
5.3. Privacy Protection
This section studies the effectiveness of three popular privacy protection techniques (DP, QT, DR) for preventing SV-driven attacks, including FIA (Luo et al., 2022), which may occur in RI tasks, and MIA (Wang et al., 2023), which may occur in DV tasks. We also study the impacts of the three techniques on SV’s effectiveness.
The implementation of two attacks and three countermeasures is based on previous papers (Luo et al., 2022; Wang et al., 2023). We follow Luo et al. (2022) to use two random distributions, Uniform(0,1) and Gaussian(0.5,$0.25^{2}$), to perform FIA (called FIAU and FIAG respectively), and follow Wang et al. (2023) to implement MIA. We tune the strength of privacy protection from low to high by varying (1) the standard deviation of noise generated by DP from 0.1, 0.5 to 0.9, (2) the number of distinctive discrete Shapley values produced by QT from $0.9n$, $0.5n$ to $0.1n$, and (3) the number of dimensions reduced for $\vec{\phi}$ by DR from $0.1n$, $0.5n$ to $0.9n$. We measure the performance in preventing SV-driven FIA by the MAE metric used by Luo et al. (2022) and measure the effectiveness of preventing SV-driven MIA by the AUROC score used by Wang et al. (2023). As both RI and DV tasks rely on the ranking of the final SVs to find the top important data features and samples, the impact of privacy protection techniques on SV effectiveness is measured by the variance in SV ranking results before and after protection.
For preventing FIA, DP is the most effective, while QT and DR need tuning the privacy protection strength to a high level for an apparent MAE enlargement. Despite this, DP renders a much larger impact on SV ranking (thus affects more on SV’s effectiveness) under the setting of low or median privacy protection strength. These phenomena are mainly because DP adds noise to SVs of all features no matter how the privacy protection strength is varied, while QT and DR alter SVs of only 10% (or 50%) data features when the strength is set to a low (or median) level.
For preventing MIA, QT and DR are effective in most cases, while DP leads to AUROC larger than the results achieved without privacy protection in some cases, e.g., DV-Bank, DV-Ttt. Moreover, different from the results of preventing FIA, enhancing the privacy protection strength may not result in better effectiveness (lower AUROC) of the three techniques in preventing MIA. For example, in DV-Wind, QT performs the worst when setting a high privacy protection strength. These outcomes stem from the inability of QT, DR, and DP to rigorously enforce indistinguishability between two Shapley value distributions: the IN distribution, computed when the target sample is included in the dataset, and the OUT distribution, computed when the target sample is excluded (Wang et al., 2023). Specifically, due to randomness, DP may inject markedly dissimilar noise into SVs used for generating IN and OUT distributions and thus enlarge the difference between the two distributions, making the success of MIA much easier. QT and DR map the original SVs into a new value space, in which the difference between SVs used for generating IN and OUT distributions might be more distinguishable, negatively affecting the efficacy in preventing SV-driven MIA.
@@ 5.4. SV Interpretations
This section examines the correctness of the mainstream SV interpretation paradigm, i.e., utility-based, in explaining SVs of different types of players defined in the current DA domain. This paradigm points out that the larger the SV of a player, the more the impacts that the player could pose on the DA task’s overall utility. Therefore, we observe the change in the overall utility of DA tasks caused by removing or adding players.
@@
Research Direction 6: In-depth investigation on key factors affecting SV interpretations. SV may not be correctly interpreted or fail to serve its desired application purposes due to two points. One is the difference between the exact SV and the approximate SV. The other one is the mismatch between the DA task and the implicit assumption of SV. For example, DV-Wind and DSV-2Dplanes tasks, as discussed above, mismatching with the assumption that the average contribution is appropriate to fairly distribute the overall utility of the task among players. Another example is the RI task in Figure 1, where pixels in an image interact mutually to influence the model predictions on the label of this image, mismatching with the assumption that the contribution of any individual player to the cooperative game is independent of its interactions with other players. Future works mining the deep understanding of these factors under more types of DA tasks to enhance the effectiveness of SV are anticipated.
Research Direction 7: Exploration on supports for complicated DA tasks. The real-world DA tasks are complicated, containing mutually dependent players (Chu and Chan, 2020; Olsen et al., 2022; Aas et al., 2021b, a; Frye et al., 2021) and real-time dynamic updates on the player set (Sun et al., 2024; Xia et al., 2025). For example, the RI task in Figure 1, where the influence of one pixel on the prediction from the image classification model tends to interrelate with nearby pixels. Besides, the dynamic real-time updates of the player set (Sun et al., 2024; Xia et al., 2025) are common in analytical tasks built upon the databases (Zhang et al., 2023b; Xia et al., 2025; Padala et al., 2025). However, most existing SV applications are ill-suited for these tasks due to the assumptions that the players are independent and the player set is static and immutable. Although conditional SV (Chen et al., 2023) and dynamic SV (Zhang et al., 2023b) have been devised for complicated tasks, these new types of SV do not treat the mutual dependency among players and their real-time dynamic updates as a single interrelated problem, thus are limited in practical applications. It is expected that future work can devise more SV applications under the complicated yet realistic DA task settings.
@@ 6. Conclusion
This paper comprehensively studied the Shapley value applied throughout the data analytics workflow. We summarized the critical variables (i.e., the player and the utility function) in designing SV applications for DA and clarified the essential functionalities of SV for data scientists. We condensed the technical challenges of applying SV in DA and discussed the related arts, qualitatively and quantitatively. The conclusions of experimental evaluations based on our development framework, SVBench, support our findings in a synthetic review. At last, we identified the limitations of current efforts and offered insights into the directions of future work.
Acknowledgements.
This work was funded by National Key Research and Development Program of China (Grant No: 2022YFB2703100), the Pioneer R&D Program of Zhejiang (No. 2024C01021), and Zhejiang Province “Leading Talent of Technological Innovation Program” (No. 2023R5214). We also appreciate anonymous reviewers for their valuable feedback and thank Jiaqi Chai for proofreading this paper.